OpenStudy (dls):

A particle executing SHM with time period of 4s. If its A=5cm.Calculate 1)Displacement 2)Velocity 3)Acceleration of the particle,1s after it starts oscillating.

OpenStudy (dls):

\[\LARGE y=asinwt\] \[\LARGE y=5\sin \frac{2 \pi}{T} \times 4\]

OpenStudy (dls):

@experimentX

OpenStudy (dls):

First I want to know the difference between small t and capital T and I think im missing something here though.

OpenStudy (experimentx):

t is your time (variable) and T is your Time Period

OpenStudy (dls):

something more about small t?

OpenStudy (experimentx):

look at your watch

OpenStudy (dls):

i dont have one :|

OpenStudy (experimentx):